OpenStudy (anonymous):

3 identical door prizes are to be given to 3 lucky people in a crowd of 100. In how many ways can this be done?

OpenStudy (lgbasallote):

when i read 3 doors i thought monty hall -__- guess not

OpenStudy (dumbcow):

use combinations....you are choosing 3 out of 100 \[\left(\begin{matrix}100 \\ 3\end{matrix}\right) = \frac{100!}{3!* 97!}\]

OpenStudy (dumbcow):

there was a monty hall problem on here the other day :)

OpenStudy (lgbasallote):

exactly :p that's how i found out about it

OpenStudy (anonymous):

ahaha oh wow thanks!

OpenStudy (dumbcow):

welcome